Charges in 1D Q2 A charge of -1.9 x 10*9 is at the origin and a charge of 6.1 x 10 C is on the x-axis at x - 3 m. At what two locations on the x-axis (Xpositive. Xnegative) is the potential zero? m Reset Enter Xpositive 0 NO m Help Reset Enter negative0 0 NO
Answer #1

-1gx1 81 k9x1o 21 6.1 3-2 21 3-거 3 2/0526 20 3.210526)2(= 3-2 4-21052674= 3 Negative ㅋ 7t 3+ 컨 (k)(y.gx1δ?) k(6./X/og) _6 3 2 10526 つし 3.210.526 ) 3t 2.210.5 2 6 )21 드 3 2t